# Mathematical Language Models: A Survey

### v3

@@ 6. Challenges and Further Directions
Mathematical LLMs reside in the phenomena of hallucination and faithfulness, where the generated output may lack factual accuracy or logical grounding (Rawte et al., 2023; Ji et al., 2023). This phenomenon leads to the production of erroneous or misleading mathematical results, undermining the reliability of the model’s outputs. Some studies explored to address this problem by integrating extra knowledge (He et al., 2022), reinforcement learning from human feedback (Lightman et al., 2023), tools (Schick et al., 2023; Parisi et al., 2022), and verify-based methods (Zhou et al., 2023d; Ling et al., 2023; Li et al., 2023b; Zhao et al., 2023a; He et al., 2022; Shridhar et al., 2023). However, the improvement of the hallucination phenomena is limited, which influences the trustworthiness and utility of mathematical language models in practical applications and scholarly pursuits.
@@
While previous models exhibit remarkable capabilities in understanding and manipulating existing mathematical concepts, their ability to autonomously devise and rigorously prove entirely new theorems presents a formidable hurdle. The development of novel mathematical theorems demands not only profound mathematical reasoning but also creative and insightful problem-solving abilities, aspects that necessitate a deeper understanding of abstract mathematical concepts beyond what the models have been trained on. Davies et al. (Davies et al., 2021) applied machine learning techniques to discover potential patterns and relationships among mathematical entities. Recently, Bernardino et al. (Bernardino Romera-Paredes et al., 2023) proposed FunSearch to find the first discoveries made for established open problems using LLMs. Bridging the gap between the models’ proficiency in assimilating existing mathematical knowledge and their capability to generate novel, verifiable, and impactful theorems represents a significant frontier in leveraging LLMs for the advancement of mathematical knowledge and discovery.
@@
The training data significantly influences the performance of language models, particularly in LLMs (Ding et al., 2023). High-quality and diverse training data can assist the model in enhancing its mathematical reasoning capabilities (Luo et al., 2023). As discussed in Section 4.1, while there have been limited studies on constructing instruction data through LLMs, these efforts have only considered building from a small set of mathematical reasoning datasets, such as GSM8k (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b). High-quality and diverse mathematical instruction data remains scarce. There is a necessity to explore and investigate additional forms and construction methods for mathematical train data. Additionally, the generation of mathematical training datasets in a multimodal context is also a promising direction.
@@ 7. Conclusions
The survey elucidates the pivotal role of mathematical LMs in reshaping the landscape of mathematical problem-solving, leveraging a spectrum of models, from pre-trained language models (PLMs) to large-scale language models (LLMs), to address diverse mathematical tasks. Our taxonomical delineation of mathematical tasks and methods provides a systematic framework for comprehending the intricacies of LMs-based methodologies, distinguishing between arithmetic calculation, mathematical reasoning, and various algorithmic approaches employed in these models. The compilation of over 60 diverse mathematical datasets, categorized meticulously into training, benchmark, and augmented datasets, underscores the pivotal role of data in advancing mathematical research, facilitating informed research endeavors within distinct mathematical contexts. Moreover, by critically addressing challenges such as faithfulness, multi-modality, uncertainty, evaluation, theorem creation, application, and Data scarcity, this survey paves the way for future investigations aimed at refining and advancing mathematical LMs capabilities. By shedding light on the current state-of-the-art, challenges, and avenues for future exploration, we envision this comprehensive overview to be a cornerstone in driving innovation and shaping the trajectory of mathematical LMs research, ultimately contributing to the evolving landscape of mathematics and artificial intelligence.
@@
  • Al-Hossami et al. (2023b) Erfan Al-Hossami, Razvan Bunescu, Ryan Teehan, Laurel Powell, Khyati Mahajan, and Mohsen Dorodchi. 2023b. Socratic questioning of novice debuggers: A benchmark dataset and preliminary evaluations. In Proceedings of the 18th Workshop on Innovative Use of NLP for Building Educational Applications (BEA 2023). 709–726.
  • Alghamdi et al. (2022) Reem Alghamdi, Zhenwen Liang, and Xiangliang Zhang. 2022. ArMATH: a Dataset for Solving Arabic Math Word Problems. In Proceedings of the Thirteenth Language Resources and Evaluation Conference. 351–362.
  • Amini et al. (2019) Aida Amini, Saadia Gabriel, Shanchuan Lin, Rik Koncel-Kedziorski, Yejin Choi, and Hannaneh Hajishirzi. 2019. MathQA: Towards Interpretable Math Word Problem Solving with Operation-Based Formalisms. In Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers). 2357–2367.
  • Ang et al. (2023) Beng Heng Ang, Sujatha Das Gollapalli, and See Kiong Ng. 2023. Socratic Question Generation: A Novel Dataset, Models, and Evaluation. In Proceedings of the 17th Conference of the European Chapter of the Association for Computational Linguistics. 147–165.
  • Antol et al. (2015) Stanislaw Antol, Aishwarya Agrawal, Jiasen Lu, Margaret Mitchell, Dhruv Batra, C Lawrence Zitnick, and Devi Parikh. 2015. Vqa: Visual question answering. In Proceedings of the IEEE international conference on computer vision. 2425–2433.
  • Austin et al. (2021) Jacob Austin, Augustus Odena, Maxwell Nye, Maarten Bosma, Henryk Michalewski, David Dohan, Ellen Jiang, Carrie Cai, Michael Terry, Quoc Le, et al. 2021. Program synthesis with large language models. arXiv preprint arXiv:2108.07732 (2021).
  • Azerbayev et al. (2023) Zhangir Azerbayev, Bartosz Piotrowski, Hailey Schoelkopf, Edward W Ayers, Dragomir Radev, and Jeremy Avigad. 2023. Proofnet: Autoformalizing and formally proving undergraduate-level mathematics. arXiv preprint arXiv:2302.12433 (2023).
@@
  • Bansal et al. (2019) Kshitij Bansal, Sarah Loos, Markus Rabe, Christian Szegedy, and Stewart Wilcox. 2019. Holist: An environment for machine learning of higher order logic theorem proving. In International Conference on Machine Learning. PMLR, 454–463.
  • Berg-Kirkpatrick and Spokoyny (2020) Taylor Berg-Kirkpatrick and Daniel Spokoyny. 2020. An empirical investigation of contextualized number prediction. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), Bonnie Webber, Trevor Cohn, Yulan He, and Yang Liu (Eds.). Association for Computational Linguistics, 4754–4764. https://doi.org/10.18653/v1/2020.emnlp-main.385
@@
  • Bertot and Castéran (2013) Yves Bertot and Pierre Castéran. 2013. Interactive theorem proving and program development: Coq’Art: the calculus of inductive constructions. Springer Science & Business Media.
  • Besta et al. (2023) Maciej Besta, Nils Blach, Ales Kubicek, Robert Gerstenberger, Lukas Gianinazzi, Joanna Gajda, Tomasz Lehmann, Michal Podstawski, Hubert Niewiadomski, Piotr Nyczyk, and Torsten Hoefler. 2023. Graph of Thoughts: Solving Elaborate Problems with Large Language Models. http://arxiv.org/abs/2308.09687 arXiv:2308.09687 [cs].
@@
  • Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al. 2020. Language models are few-shot learners. Advances in neural information processing systems 33 (2020), 1877–1901.
  • Chang (2023) Edward Y Chang. 2023. Prompting large language models with the socratic method. In 2023 IEEE 13th Annual Computing and Communication Workshop and Conference (CCWC). IEEE, 0351–0360.
  • Chen et al. (2022a) Wenhu Chen, Xueguang Ma, Xinyi Wang, and William W Cohen. 2022a. Program of thoughts prompting: Disentangling computation from reasoning for numerical reasoning tasks. arXiv preprint arXiv:2211.12588 (2022).
  • Chen et al. (2022b) Yanda Chen, Ruiqi Zhong, Sheng Zha, George Karypis, and He He. 2022b. Meta-learning via Language Model In-context Tuning. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). 719–730.
  • Chen et al. (2021) Zhiyu Chen, Wenhu Chen, Charese Smiley, Sameena Shah, Iana Borova, Dylan Langdon, Reema Moussa, Matt Beane, Ting-Hao Huang, Bryan R Routledge, et al. 2021. FinQA: A Dataset of Numerical Reasoning over Financial Data. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing. 3697–3711.
  • Chowdhery et al. (2022) Aakanksha Chowdhery, Sharan Narang, Jacob Devlin, Maarten Bosma, Gaurav Mishra, Adam Roberts, Paul Barham, Hyung Won Chung, Charles Sutton, Sebastian Gehrmann, et al. 2022. Palm: Scaling language modeling with pathways. arXiv preprint arXiv:2204.02311 (2022).
@@
  • Chung et al. (2022) Hyung Won Chung, Le Hou, Shayne Longpre, Barret Zoph, Yi Tay, William Fedus, Yunxuan Li, Xuezhi Wang, Mostafa Dehghani, Siddhartha Brahma, et al. 2022. Scaling instruction-finetuned language models. arXiv preprint arXiv:2210.11416 (2022).
  • Clark et al. (2021) Peter Clark, Oren Etzioni, Daniel Khashabi, Tushar Khot, Bhavana Dalvi Mishra, Kyle Richardson, Ashish Sabharwal, Carissa Schoenick, Oyvind Tafjord, Niket Tandon, Sumithra Bhakthavatsalam, Dirk Groeneveld, Michal Guerquin, and Michael Schmitz. 2021. From ’F’ to ’a’ on the N.Y. Regents Science Exams: An Overview of the Aristo Project. arXiv:1909.01958 [cs]
  • Cobbe et al. (2021) Karl Cobbe, Vineet Kosaraju, Mohammad Bavarian, Mark Chen, Heewoo Jun, Lukasz Kaiser, Matthias Plappert, Jerry Tworek, Jacob Hilton, Reiichiro Nakano, Christopher Hesse, and John Schulman. 2021. Training Verifiers to Solve Math Word Problems. arXiv:2110.14168 [cs.LG]
@@
  • Dagan et al. (2023) Gautier Dagan, Frank Keller, and Alex Lascarides. 2023. Dynamic Planning with a LLM. arXiv preprint arXiv:2308.06391 (2023).
@@
  • Devlin et al. (2019) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019. BERT: Pre-training of Deep Bidirectional Transformers for Language Understanding. arXiv:1810.04805 [cs]
  • Ding et al. (2023) Ning Ding, Yulin Chen, Bokai Xu, Yujia Qin, Zhi Zheng, Shengding Hu, Zhiyuan Liu, Maosong Sun, and Bowen Zhou. 2023. Enhancing Chat Language Models by Scaling High-quality Instructional Conversations. http://arxiv.org/abs/2305.14233 arXiv:2305.14233 [cs].
  • Drori et al. (2022) Iddo Drori, Sarah Zhang, Reece Shuttleworth, Leonard Tang, Albert Lu, Elizabeth Ke, Kevin Liu, Linda Chen, Sunny Tran, Newman Cheng, et al. 2022. A neural network solves, explains, and generates university math problems by program synthesis and few-shot learning at human level. Proceedings of the National Academy of Sciences 119, 32 (2022), e2123433119.
  • Dua et al. (2019) Dheeru Dua, Yizhong Wang, Pradeep Dasigi, Gabriel Stanovsky, Sameer Singh, and Matt Gardner. 2019. DROP: A Reading Comprehension Benchmark Requiring Discrete Reasoning Over Paragraphs. In Proceedings of NAACL-HLT. 2368–2378.
  • Duan et al. (2023) Jinhao Duan, Hao Cheng, Shiqi Wang, Chenan Wang, Alex Zavalny, Renjing Xu, Bhavya Kailkhura, and Kaidi Xu. 2023. Shifting attention to relevance: Towards the uncertainty estimation of large language models. arXiv preprint arXiv:2307.01379 (2023).
  • Elazar et al. (2019) Yanai Elazar, Abhijit Mahabal, Deepak Ramachandran, Tania Bedrax-Weiss, and Dan Roth. 2019. How Large Are Lions? Inducing Distributions over Quantitative Attributes. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics. 3973–3983.
  • Feigenbaum et al. (1963) Edward A Feigenbaum, Julian Feldman, et al. 1963. Computers and thought. Vol. 7. New York McGraw-Hill.
  • Feng et al. (2022) Yu Feng, Jing Zhang, Xiaokang Zhang, Lemao Liu, Cuiping Li, and Hong Chen. 2022. Injecting Numerical Reasoning Skills into Knowledge Base Question Answering Models. arXiv:2112.06109 [cs]
  • First et al. (2023) Emily First, Markus N Rabe, Talia Ringer, and Yuriy Brun. 2023. Baldur: whole-proof generation and repair with large language models. arXiv preprint arXiv:2303.04910 (2023).
@@
  • Forbes and Choi (2017) Maxwell Forbes and Yejin Choi. 2017. Verb Physics: Relative Physical Knowledge of Actions and Objects. In Proceedings of the 55th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). 266–276.
  • Fu et al. (2023) Yao Fu, Hao Peng, Ashish Sabharwal, Peter Clark, and Tushar Khot. 2023. Complexity-Based Prompting for Multi-step Reasoning. In The Eleventh International Conference on Learning Representations.
  • Gao et al. (2023) Luyu Gao, Aman Madaan, Shuyan Zhou, Uri Alon, Pengfei Liu, Yiming Yang, Jamie Callan, and Graham Neubig. 2023. Pal: Program-aided language models. In International Conference on Machine Learning. PMLR, 10764–10799.
@@
  • Gelernter et al. (1960) Herbert Gelernter, James R Hansen, and Donald W Loveland. 1960. Empirical explorations of the geometry theorem machine. In Papers presented at the May 3-5, 1960, western joint IRE-AIEE-ACM computer conference. 143–149.
  • Geva et al. (2020) Mor Geva, Ankit Gupta, and Jonathan Berant. 2020. Injecting Numerical Reasoning Skills into Language Models. arXiv:2004.04487 [cs]
  • Gonthier et al. (2013) Georges Gonthier, Andrea Asperti, Jeremy Avigad, Yves Bertot, Cyril Cohen, François Garillot, Stéphane Le Roux, Assia Mahboubi, Russell O’Connor, Sidi Ould Biha, et al. 2013. A machine-checked proof of the odd order theorem. In International conference on interactive theorem proving. Springer, 163–179.
  • Gou et al. (2023a) Zhibin Gou, Zhihong Shao, Yeyun Gong, Yelong Shen, Yujiu Yang, Nan Duan, and Weizhu Chen. 2023a. CRITIC: Large Language Models Can Self-Correct with Tool-Interactive Critiquing. arXiv:2305.11738 [cs.CL]
  • Gou et al. (2023b) Zhibin Gou, Zhihong Shao, Yeyun Gong, Yujiu Yang, Minlie Huang, Nan Duan, Weizhu Chen, et al. 2023b. Tora: A tool-integrated reasoning agent for mathematical problem solving. arXiv preprint arXiv:2309.17452 (2023).

### v4

@@ 7. Challenges and Further Directions
Mathematical LLMs reside in the phenomena of hallucination and faithfulness, where the generated output may lack factual accuracy or logical grounding (Rawte et al., 2023; Ji et al., 2023). This phenomenon leads to the production of erroneous or misleading mathematical results, undermining the reliability of the model’s outputs. Some studies explored to address this problem by integrating extra knowledge (He et al., 2022), reinforcement learning from human feedback (Lightman et al., 2024), tools (Schick et al., 2023; Parisi et al., 2022), and verify-based methods (Zhou et al., 2023d; Ling et al., 2023; Li et al., 2023b; Zhao et al., 2023a; He et al., 2022; Shridhar et al., 2023). However, the improvement of the hallucination phenomena is limited, which influences the trustworthiness and utility of mathematical language models in practical applications and scholarly pursuits.
@@
While previous models exhibit remarkable capabilities in understanding and manipulating existing mathematical concepts, their ability to autonomously devise and rigorously prove entirely new theorems presents a formidable hurdle. The development of novel mathematical theorems demands not only profound mathematical reasoning but also creative and insightful problem-solving abilities, aspects that necessitate a deeper understanding of abstract mathematical concepts beyond what the models have been trained on. Davies et al. (Davies et al., 2021) applied machine learning techniques to discover potential patterns and relationships among mathematical entities. Recently, Bernardino et al. (Bernardino Romera-Paredes et al., 2023) proposed FunSearch to find the first discoveries made for established open problems using LLMs. Bridging the gap between the models’ proficiency in assimilating existing mathematical knowledge and their capability to generate novel and impactful theorems represents a significant frontier in leveraging LLMs for the advancement of mathematical discovery.
@@
The training data significantly influences the performance of language models, particularly in LLMs (Ding et al., 2023). High-quality and diverse training data can assist the model in enhancing its mathematical reasoning capabilities (Luo et al., 2023). As discussed in Section 4.1, while there have been limited studies on constructing instruction data through LLMs, these efforts have only considered building from a small set of mathematical reasoning datasets, such as GSM8k (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b). High-quality and diverse mathematical instruction data remains scarce. It is necessary to explore additional forms and construction methods for mathematical train data. Additionally, the generation of mathematical training datasets in a multimodal context is also a promising direction.
@@ 8. Conclusions
The survey elucidates the pivotal role of mathematical LMs in reshaping the landscape of mathematical problem-solving, leveraging a spectrum of models, from PLMs to LLMs, to address diverse mathematical tasks. Our taxonomical delineation of mathematical tasks and methods provides a systematic framework for comprehending the intricacies of LMs-based methodologies, distinguishing between arithmetic calculation, mathematical reasoning, and various algorithmic approaches employed in these models. We also outline the features and effectiveness of current mathematical LMs to gain a deeper understanding of these techniques. We gather over 60 diverse mathematical datasets, carefully organized into training, benchmark, and augmented categories, highlighting how crucial data is in pushing forward mathematical research and supporting insightful inquiry in various mathematical fields. Additionally, by thoughtfully tackling challenges like faithfulness, multi-modality, uncertainty, evaluation, theorem creation, application, and data scarcity, this survey opens doors for future studies aimed at enhancing the capabilities of mathematical LMs. By exploring the current advancements, obstacles, and paths for future research, we see this thorough overview as a key resource for fostering innovation and guiding the future of mathematical LMs research, all while contributing to the exciting developments in mathematics and artificial intelligence.
Acknowledgements.
The authors wish to thank the reviewers for their helpful comments and suggestions. This research is funded by the National Science and Technology Major Project (No. 2021ZD0114002), the National Nature Science Foundation of China (No. 62477010, No.62307028), the Science and Technology Commission of Shanghai Municipality Grant (No. 22511105901, No. 21511100402), Shanghai Science and Technology Innovation Action Plan (No. 24YF2710100) and Shanghai Special Project to Promote High-quality Industrial Development (No. RZ-CYAI-01-24-0288).
@@ Acknowledgements.
  • Al-Hossami et al. (2023b) Erfan Al-Hossami, Razvan Bunescu, Ryan Teehan, Laurel Powell, Khyati Mahajan, and Mohsen Dorodchi. 2023b. Socratic questioning of novice debuggers: A benchmark dataset and preliminary evaluations. In BEA. 709–726.
  • Alghamdi et al. (2022) Reem Alghamdi, Zhenwen Liang, and Xiangliang Zhang. 2022. ArMATH: a Dataset for Solving Arabic Math Word Problems. In LREC. 351–362.
  • Amini et al. (2019) Aida Amini, Saadia Gabriel, Shanchuan Lin, Rik Koncel-Kedziorski, Yejin Choi, and Hannaneh Hajishirzi. 2019. MathQA: Towards Interpretable Math Word Problem Solving with Operation-Based Formalisms. In NAACL. 2357–2367.
  • Ang et al. (2023) Beng Heng Ang, Sujatha Das Gollapalli, and See Kiong Ng. 2023. Socratic Question Generation: A Novel Dataset, Models, and Evaluation. In EACL. 147–165.
  • Anonymous (2024) Anonymous. 2024. Diving into Self-Evolve Training for Multimodal Reasoning. In Submitted to The Thirteenth International Conference on Learning Representations.
  • Antol et al. (2015) Stanislaw Antol, Aishwarya Agrawal, Jiasen Lu, Margaret Mitchell, Dhruv Batra, C Lawrence Zitnick, and Devi Parikh. 2015. Vqa: Visual question answering. In ICCV. 2425–2433.
  • Austin et al. (2021) Jacob Austin, Augustus Odena, Maxwell Nye, Maarten Bosma, Henryk Michalewski, David Dohan, Ellen Jiang, Carrie Cai, Michael Terry, Quoc Le, et al. 2021. Program synthesis with large language models. arXiv (2021).
  • Azerbayev et al. (2023) Zhangir Azerbayev, Bartosz Piotrowski, Hailey Schoelkopf, Edward W Ayers, Dragomir Radev, and Jeremy Avigad. 2023. Proofnet: Autoformalizing and formally proving undergraduate-level mathematics. arXiv (2023).
@@
  • Bansal et al. (2019) Kshitij Bansal, Sarah Loos, Markus Rabe, Christian Szegedy, and Stewart Wilcox. 2019. Holist: An environment for machine learning of higher order logic theorem proving. In ICML. 454–463.
  • Berg-Kirkpatrick and Spokoyny (2020) Taylor Berg-Kirkpatrick and Daniel Spokoyny. 2020. An empirical investigation of contextualized number prediction. In EMNLP. 4754–4764.
@@
  • Bertot and Castéran (2013) Yves Bertot and Pierre Castéran. 2013. Interactive theorem proving and program development: Coq’Art: the calculus of inductive constructions.
  • Besta et al. (2023) Maciej Besta, Nils Blach, Ales Kubicek, Robert Gerstenberger, Lukas Gianinazzi, Joanna Gajda, Tomasz Lehmann, Michal Podstawski, Hubert Niewiadomski, Piotr Nyczyk, and Torsten Hoefler. 2023. Graph of Thoughts: Solving Elaborate Problems with Large Language Models.
@@
  • Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al. 2020. Language models are few-shot learners. NeurIPS 33 (2020), 1877–1901.
  • Chang (2023) Edward Y Chang. 2023. Prompting large language models with the socratic method. In CCWC. 0351–0360.
  • Chaslot et al. (2008) Guillaume Chaslot, Sander Bakkes, Istvan Szita, and Pieter Spronck. 2008. Monte-carlo tree search: A new framework for game ai. In AAAI, Vol. 4. 216–217.
  • Chen et al. (2024a) Huayu Chen, Guande He, Lifan Yuan, Ganqu Cui, Hang Su, and Jun Zhu. 2024a. Noise contrastive alignment of language models with explicit rewards. arXiv (2024).
  • Chen et al. (2021b) Jiaqi Chen, Jianheng Tang, Jinghui Qin, Xiaodan Liang, Lingbo Liu, Eric P Xing, and Liang Lin. 2021b. GeoQA: A geometric question answering benchmark towards multimodal numerical reasoning. arXiv (2021).
  • Chen et al. (2024c) Nuo Chen, Ning Wu, Jianhui Chang, and Jia Li. 2024c. ControlMath: Controllable Data Generation Promotes Math Generalist Models. arXiv (2024).
  • Chen et al. (2024b) Qiguang Chen, Libo Qin, Jin Zhang, Zhi Chen, Xiao Xu, and Wanxiang Che. 2024b. M3CoT: A Novel Benchmark for Multi-Domain Multi-step Multi-modal Chain-of-Thought. In ACL. 8199–8221.
  • Chen et al. (2022a) Wenhu Chen, Xueguang Ma, Xinyi Wang, and William W Cohen. 2022a. Program of thoughts prompting: Disentangling computation from reasoning for numerical reasoning tasks. arXiv (2022).
  • Chen et al. (2022b) Yanda Chen, Ruiqi Zhong, Sheng Zha, George Karypis, and He He. 2022b. Meta-learning via Language Model In-context Tuning. In ACL. 719–730.
  • Chen et al. (2021a) Zhiyu Chen, Wenhu Chen, Charese Smiley, Sameena Shah, Iana Borova, Dylan Langdon, Reema Moussa, Matt Beane, Ting-Hao Huang, Bryan R Routledge, et al. 2021a. FinQA: A Dataset of Numerical Reasoning over Financial Data. In EMNLP. 3697–3711.
  • Chen et al. (2024d) Zhe Chen, Jiannan Wu, Wenhai Wang, Weijie Su, Guo Chen, Sen Xing, Muyan Zhong, Qinglong Zhang, Xizhou Zhu, Lewei Lu, et al. 2024d. Internvl: Scaling up vision foundation models and aligning for generic visual-linguistic tasks. In CVPR. 24185–24198.
  • Chernyshev et al. (2024) Konstantin Chernyshev, Vitaliy Polshkov, Ekaterina Artemova, Alex Myasnikov, Vlad Stepanov, Alexei Miasnikov, and Sergei Tilga. 2024. U-MATH: A University-Level Benchmark for Evaluating Mathematical Skills in LLMs. arXiv:2412.03205
  • Chowdhery et al. (2022) Aakanksha Chowdhery, Sharan Narang, Jacob Devlin, Maarten Bosma, Gaurav Mishra, Adam Roberts, Paul Barham, Hyung Won Chung, Charles Sutton, Sebastian Gehrmann, et al. 2022. Palm: Scaling language modeling with pathways. arXiv (2022).
@@
  • Chung et al. (2022) Hyung Won Chung, Le Hou, Shayne Longpre, Barret Zoph, Yi Tay, William Fedus, Yunxuan Li, Xuezhi Wang, Mostafa Dehghani, Siddhartha Brahma, et al. 2022. Scaling instruction-finetuned language models. arXiv (2022).
  • Clark et al. (2021) Peter Clark, Oren Etzioni, Daniel Khashabi, Tushar Khot, Bhavana Dalvi Mishra, Kyle Richardson, Ashish Sabharwal, Carissa Schoenick, Oyvind Tafjord, Niket Tandon, Sumithra Bhakthavatsalam, Dirk Groeneveld, Michal Guerquin, and Michael Schmitz. 2021. From ’F’ to ’a’ on the N.Y. Regents Science Exams: An Overview of the Aristo Project. arXiv:1909.01958
  • Cobbe et al. (2021) Karl Cobbe, Vineet Kosaraju, Mohammad Bavarian, Mark Chen, Heewoo Jun, Lukasz Kaiser, Matthias Plappert, Jerry Tworek, Jacob Hilton, Reiichiro Nakano, Christopher Hesse, and John Schulman. 2021. Training Verifiers to Solve Math Word Problems. arXiv:2110.14168
@@
  • Dagan et al. (2023) Gautier Dagan, Frank Keller, and Alex Lascarides. 2023. Dynamic Planning with a LLM. arXiv (2023).
@@
  • Deng and Mineiro (2024) Yihe Deng and Paul Mineiro. 2024. Flow-DPO: Improving LLM Mathematical Reasoning through Online Multi-Agent Learning. arXiv (2024).
  • Devlin et al. (2019) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019. BERT: Pre-training of Deep Bidirectional Transformers for Language Understanding. arXiv:1810.04805
  • Ding et al. (2023) Ning Ding, Yulin Chen, Bokai Xu, Yujia Qin, Shengding Hu, Zhiyuan Liu, Maosong Sun, and Bowen Zhou. 2023. Enhancing Chat Language Models by Scaling High-quality Instructional Conversations. In EMNLP. 3029–3051.
  • Ding et al. (2024) Yuyang Ding, Hanglei Hu, Jie Zhou, Qin Chen, Bo Jiang, and Liang He. 2024. Boosting Large Language Models with Socratic Method for Conversational Mathematics Teaching. In CIKM. 3730–3735.
  • Drori et al. (2022) Iddo Drori, Sarah Zhang, Reece Shuttleworth, Leonard Tang, Albert Lu, Elizabeth Ke, Kevin Liu, Linda Chen, Sunny Tran, Newman Cheng, et al. 2022. A neural network solves, explains, and generates university math problems by program synthesis and few-shot learning at human level. PNAS 119, 32 (2022), e2123433119.
  • Dua et al. (2019) Dheeru Dua, Yizhong Wang, Pradeep Dasigi, Gabriel Stanovsky, Sameer Singh, and Matt Gardner. 2019. DROP: A Reading Comprehension Benchmark Requiring Discrete Reasoning Over Paragraphs. In NAACL. 2368–2378.
  • Duan et al. (2023) Jinhao Duan, Hao Cheng, Shiqi Wang, Chenan Wang, Alex Zavalny, Renjing Xu, Bhavya Kailkhura, and Kaidi Xu. 2023. Shifting attention to relevance: Towards the uncertainty estimation of large language models. arXiv (2023).
  • Elazar et al. (2019) Yanai Elazar, Abhijit Mahabal, Deepak Ramachandran, Tania Bedrax-Weiss, and Dan Roth. 2019. How Large Are Lions? Inducing Distributions over Quantitative Attributes. In ACL. 3973–3983.
  • Ethayarajh et al. (2024) Kawin Ethayarajh, Winnie Xu, Niklas Muennighoff, Dan Jurafsky, and Douwe Kiela. 2024. Kto: Model alignment as prospect theoretic optimization. arXiv (2024).
  • Feigenbaum et al. (1963) Edward A Feigenbaum, Julian Feldman, et al. 1963. Computers and thought. Vol. 7.
  • Feng et al. (2022) Yu Feng, Jing Zhang, Xiaokang Zhang, Lemao Liu, Cuiping Li, and Hong Chen. 2022. Injecting Numerical Reasoning Skills into Knowledge Base Question Answering Models. arXiv:2112.06109
  • First et al. (2023) Emily First, Markus N Rabe, Talia Ringer, and Yuriy Brun. 2023. Baldur: whole-proof generation and repair with large language models. arXiv (2023).
@@
  • Forbes and Choi (2017) Maxwell Forbes and Yejin Choi. 2017. Verb Physics: Relative Physical Knowledge of Actions and Objects. In ACL. 266–276.
  • Fu et al. (2023) Yao Fu, Hao Peng, Ashish Sabharwal, Peter Clark, and Tushar Khot. 2023. Complexity-Based Prompting for Multi-step Reasoning. In ICLR.
  • Gao et al. (2023) Luyu Gao, Aman Madaan, Shuyan Zhou, Uri Alon, Pengfei Liu, Yiming Yang, Jamie Callan, and Graham Neubig. 2023. Pal: Program-aided language models. In ICML. 10764–10799.
@@
  • Gelernter et al. (1960) Herbert Gelernter, James R Hansen, and Donald W Loveland. 1960. Empirical explorations of the geometry theorem machine. In western joint IRE-AIEE-ACM computer conference. 143–149.
  • Geva et al. (2020) Mor Geva, Ankit Gupta, and Jonathan Berant. 2020. Injecting Numerical Reasoning Skills into Language Models. In ACL, Dan Jurafsky, Joyce Chai, Natalie Schluter, and Joel Tetreault (Eds.). 946–958.
  • GLM et al. (2024) Team GLM, Aohan Zeng, Bin Xu, Bowen Wang, Chenhui Zhang, Da Yin, Dan Zhang, Diego Rojas, Guanyu Feng, Hanlin Zhao, et al. 2024. Chatglm: A family of large language models from glm-130b to glm-4 all tools. arXiv (2024).
  • Gonthier et al. (2013) Georges Gonthier, Andrea Asperti, Jeremy Avigad, Yves Bertot, Cyril Cohen, François Garillot, Stéphane Le Roux, Assia Mahboubi, Russell O’Connor, Sidi Ould Biha, et al. 2013. A machine-checked proof of the odd order theorem. In ITP. 163–179.
  • Gou et al. (2023a) Zhibin Gou, Zhihong Shao, Yeyun Gong, Yelong Shen, Yujiu Yang, Nan Duan, and Weizhu Chen. 2023a. CRITIC: Large Language Models Can Self-Correct with Tool-Interactive Critiquing. arXiv:2305.11738
  • Gou et al. (2023b) Zhibin Gou, Zhihong Shao, Yeyun Gong, Yujiu Yang, Minlie Huang, Nan Duan, Weizhu Chen, et al. 2023b. Tora: A tool-integrated reasoning agent for mathematical problem solving. arXiv (2023).